# ULSR-GS: Ultra Large-scale Surface Reconstruction Gaussian Splatting with Multi-View Geometric Consistency

### v1

@@ Abstract
While Gaussian Splatting (GS) demonstrates efficient and high-quality scene rendering and small area surface extraction ability, it falls short in handling large-scale aerial image surface extraction tasks. To overcome this, we present ULSR-GS, a framework dedicated to high-fidelity surface extraction in ultra-large-scale scenes, addressing the limitations of existing GS-based mesh extraction methods. Specifically, we propose a point-to-photo partitioning approach combined with a multi-view optimal view matching principle to select the best training images for each sub-region. Additionally, during training, ULSR-GS employs a densification strategy based on multi-view geometric consistency to enhance surface extraction details. Experimental results demonstrate that ULSR-GS outperforms other state-of-the-art GS-based works on large-scale aerial photogrammetry benchmark datasets, significantly improving surface extraction accuracy in complex urban environments. Project page: https://ulsrgs.github.io.
Figure 1: We present ULSR-GS, a large-scale high-fidelity surface extraction framework based on Gaussian Splatting. The three sub-regions shown in the figure are stitched together after parallel training by three GPUs on the Hospital scene of the UrbanScene3D [21] dataset.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) has emerged as a groundbreaking approach in 3D surface reconstruction and rendering, enabling high-quality scene representations and supporting the extraction of detail-rich meshes [12]. However, existing GS-based research has primarily focused on small-scale surface reconstructions [10, 11, 49, 3, 18] or large-scale scene rendering [19, 48, 22] without optimizing for large-scale urban surface reconstruction. Several factors contribute to the scarcity of GS-based applications in large-scale urban surface reconstruction: (1) Computational complexity is a significant hurdle [24, 19], as processing vast urban datasets on a single GPU demands extensive computational resources and memory, challenging the scalability of GS methods. (2) Existing large-scale GS-based methods are typically optimized for scene rendering, and their partition strategies lack the necessary adaptations for handling the mesh reconstruction tasks. (3) Reconstruction quality challenges arise due to insufficient densification processes in certain areas, especially when handling thousands of aerial oblique images [7, 8].
To tackle the challenges of GS-based large-scale surface reconstruction, we propose ULSR-GS, a large-scale surface reconstruction method combined with the point-to-photo partition method and multi-view constrained densification. Specifically, unlike methods that focus on image position-based partitioning [19, 48], we select the optimal set of views based on the matched multi-view images [42, 45, 39, 20, 36, 41], considering factors such as camera angle, proximity, and pairing quality for each point. This targeted image selection not only ensures high-quality surface reconstruction by leveraging the most informative images but also reduces redundant data processing. We also introduce a multi-view training strategy that significantly enhances the reconstruction quality. During each training iteration, we impose additional constraints based on the best-matched views corresponding to the current image being processed [42, 18]. This training strategy enables more robust and consistent reconstructions by leveraging the most informative images from multiple angles. We further incorporate this multi-view consistency constraint in the densification process. This ensures that the densification step adheres to consistent geometric relationships across views, leading to more accurate surface details.
@@
  • We present ULSR-GS, a novel method specifically designed to overcome the limitations of existing GS-based works in large-scale surface reconstruction.
  • We propose an innovative point-to-photo scene partition strategy for GS in large-scale scene mesh extraction. This method select images for each sub-region based on the best-matched views, improving reconstruction accuracy.
  • We propose multi-view consistency constraints for the additional densification process for generating detailed and accurate reconstructions in large-scale urban environments.
@@ 2.1 Large-Scale Novel View Synthesis
3D reconstruction has evolved from traditional Multi-View Stereo (MVS) [29, 41, 2] to advanced neural methods [34, 27, 44], and mesh extraction methods transform implicit representations into explicit 3D models. Poisson Surface-based methods extract surfaces from point clouds but often produce overly smooth results, lacking fine details [10]. Marching Cubes converts volumetric data into polygonal meshes, but its grid-based nature can lead to artifacts and increased memory usage [23, 18]. Neural implicit representations encode 3D geometry continuously, enabling smooth modeling and handling complex topologies [34, 44, 43]. However, these methods require significant computational resources and often struggle with fine geometric details. Mesh extraction from these representations also suffers from discretization artifacts. Neural Radiance Fields (NeRF) use volumetric rendering to achieve high-quality view synthesis [26]. Despite their impressive results, NeRF-based methods are computationally intensive, making them less practical for real-time applications [1, 32]. Recent works extract explicit geometry from NeRF, but converting density fields to surfaces in large-scale scenes remains challenging and may produce noisy results [17].
@@
Figure 2: Point-based partition strategy of ULSR-GS. (I) We partition the training scene based on the positions of the point cloud projected onto the xz-plane. (II) For each sub-scene: (a) we first select all matching images within the region; (b) based on the camera relationships of the sub-scene, we select the corresponding images for each point; (c) following the principle of multi-view optimal matching [42], we filter out images with poor view angles; and finally, (d) considering the distance relationship, we further remove images that are too distant. (III) The final reconstructed sub-regions significantly reduce the number of training images.
@@ 3.1 Multi-View Pair Selection
here, $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ are the camera centers of views $i$ and $j$, respectively. The function $G$ is a piecewise Gaussian function that favors a certain baseline angle $\theta_{0}$:
@@ 3.2 Depth Calculation in 2D Gaussian Splatting
$D_{mean}=\sum_{i}{\omega_{i}D_{i}}/(\sum_{i}{\omega_{i}D_{i}}+\epsilon)$ (3)
@@
Maximum visibility. This method focuses on selecting the depth of the most ”visible” Gaussian along the viewing ray. The depth $D_{median}$ is determined by the first Gaussian that contributes significantly to the cumulative alpha:
@@
$D_{mean}$ facilitates smoother depth calculations, whereas $D_{median}$ more effectively captures rich geometric details. however, it is highly susceptible to floating-point errors that can impact depth computation. In this paper, we adopt $D_{mean}$ as the depth representation for subsequent analyses.
@@ 4 Method
We introduce the proposed point partition method with multi-view image selection in Section 4.1. Section 4.2 introduces a multi-view densification method with depth aggregation. In Section 4.3, we introduce the loss function for the multi-view training process.
@@ 4.1 Multi-View Optimized Point Partitioning
Different from previous studies that use drone photo locations for region partitioning [19, 48, 22], our scene partitioning method is based on the initial point cloud of the scene and selects the best training images for the point cloud in each sub-region. The advantage of this method is that we can determine the boundaries of the mesh to be extracted for each sub-area at the early stage, without the need to merge the full scene.
Density-controlled boundary refinement. To effectively partition the initial points into $n\times n$ sub-regions, it is crucial to isolate the primary structural components of the scene and eliminate sparse and noisy SfM points that could distort the boundary definition.
@@
Initial view selection. As shown in Figure 2 (I), the input COLMAP SfM [29] point cloud is divided into $n\times n$ grids after density filtering. Each point in the sub-region is the feature of the images in which it was detected and matched. We first select all matched images as a coarse view selection (see Figure 2 (e)).
Source view selection. To further refine the view selection in Eq. (1), a region constraint is applied based on the camera pair distance $d_{ij}$. Only image pairs with a distance below a specified maximum threshold $D_{\max}$ are considered for matching. The final matching score $S_{ij}$:
@@
Per-point optimal view selection. Our goal is to ensure that each SfM point within a sub-region is associated with its most informative and geometrically robust image pairs. First, we project each point $P_{k}$ onto the 2D image planes of each reference image ${I}_{\text{ref}_{i}}$ and corresponding source images ${I}_{\text{src}_{i}^{1}}$, ${I}_{\text{src}_{i}^{2}}$, ${I}_{\text{src}_{i}^{3}}$ that can observe $P_{k}$. We then calculate the average euclidean distance $D_{avg}(P_{k})$ between each projected point and the corresponding centers. Among all possible four-image groups that observe $P_{k}$, we select the group with the smallest $D_{avg}(P_{k})_{min}$. This ensures that $P_{k}$ is primarily reconstructed using images where it is closest to the image centers in their respective 2D planes, thereby enhancing the reliability of its triangulation.
@@ 4.2 Adaptive Multi-View Densification
Prior works [9, 18, 7, 47] have established that finer Gaussian primitives are critical for capturing high-frequency geometric details. In ULSR-GS, we perform additional densification via a MVS-like approach [8], to address the limitations of overly smooth meshes resulting from $D_{mean}$ computations from TSDF fusion. This method projects the $D_{mean}$ into 3D space [42, 8] and combines with the RGB information of the GT image to enrich the Gaussian primitives.
Multi-view depth aggregation. In our approach, we perform weighted averaging of depth information from multiple source views. This weighting assigns a confidence score to each depth estimate based on its geometric consistency [42], ensuring that source views with higher consistency have a greater influence on the final depth estimation.
Scince we have $3$ source views from Eq. (8), with each rendered depth $D_{mean}$ denoted as $\mathbf{D}_{\text{src}_{i}}$ for $i=1,2,3$, and the depth map of the reference view denoted as $\mathbf{D}_{\text{ref}}$. For each pixel $p_{\text{ref}}$ in the reference view, the final fused depth estimate $\mathbf{D}_{\text{final}}(p_{\text{ref}})$ is obtained through weighted fusion of the source views:
@@
Adaptive depth densification. Directly projecting all depth information after checking the geometric consistency introduces excessive redundant information into the training scene [18]. This redundancy negatively impacts both the training speed and the accuracy of the scene representation. To address this issue, we introduce an adaptive densification window mask that confines the densification area. This approach eliminates inaccurate values at the edges of the depth map and adaptively accounts for non-uniform depth changes caused by varying viewpoint poses.

### v2

# inst4] organization=IPTC, ETSI de Telecomunicación,addressline=Universidad Politécnica de Madrid, city=Madrid, postcode=28040, country=Spain

@@ Abstract
Recent advances in 2D Gaussian Splatting (2DGS) have demonstrated compelling rendering efficiency and mesh extraction capabilities. However, its application to large-scale aerial photogrammetry, especially using oblique UAV imagery, remains limited due to three primary challenges: (1) suboptimal image selection in scene partitioning strategies failing to scale effectively; (2) densification pipelines that rely primarily on single-view constraints resulting in over-smoothed reconstructions and loss of fine geometric detail; and (3) the absence of multi-view geometric consistency constraints leading to surface artifacts and inconsistencies. To address these limitations, we propose ULSR-GS, a novel method tailored for high-resolution surface reconstruction in urban-scale environments. Firstly, we propose a point-to-photo partitioning strategy that segments the scene based on the sparse SfM point cloud and assigns only the most relevant images to each sub-region, which resolves key scalability bottlenecks. Secondly, we propose a multi-view guided densification strategy that enforces adaptive geometric consistency across views, overcoming the limitations of single-view-based densifications. Lastly, we introduce consistency-aware loss functions that explicitly regulate depth and normal alignment across views, significantly enhancing surface fidelity. Extensive experiments on large-scale aerial benchmark datasets demonstrate that ULSR-GS consistently outperforms existing single- and multi-GPU Gaussian Splatting methods. Furthermore, compared to MVS pipelines, our approach achieves comparable or superior geometric quality while being substantially more time-efficient, making it a practical solution for scalable 3D modeling in digital twin and urban mapping applications. Project page: https://ulsrgs.github.io.
keywords:
[…]
@@ 1 Introduction
High-fidelity 3D surface reconstruction at urban scale remains a longstanding challenge in photogrammetry and computer graphics, with critical applications in urban planning, simulation, and digital twins [51, 42, 12]. Traditionally, urban-scale surface reconstruction from aerial oblique imagery has relied heavily on Multi-View Stereo (MVS) techniques [44, 24, 56, 57, 3, 59, 71, 41, 32, 60, 14, 31, 70, 28], which reconstruct dense point clouds from image correspondences and subsequently extract meshes through surface reconstruction algorithms, such as Poisson Surface Reconstruction [17, 18] or Delaunay triangulation-based methods [23]. While effective, these traditional pipelines often encounter computational inefficiencies when scaling to urban-scale datasets comprising thousands of high-resolution images. Recent advancements in novel view synthesis (NVS) have led to from Neural Radiance Fields (NeRFs) [58, 69, 4, 1, 2] to the emergence of 3D Gaussian Splatting (3DGS) [19], a breakthrough approach enabling fast photorealistic rendering quality with unprecedented efficiency by directly optimizing sparse 3D Gaussian primitives. In addition to achieving compelling visual realism, recent studies have demonstrated that 3DGS and its variants can serve as effective tools for extracting geometry-rich meshes [16, 5, 27].
Scaling Gaussian Splatting (GS) to city-scale surface reconstruction tasks is both an engineering and algorithmic challenge. Issue (1): The vanilla GS training pipeline [19, 16, 64, 66] does not inherently support multi-GPU parallel processing. For instance, even high-performance GPUs like the RTX 4090 (24GB VRAM) can manage only approximately 8 million primitives, limiting applications to relatively small datasets such as the Mip-NeRF 360 dataset [2]. Consequently, current large-scale GS approaches predominantly rely on a divide-and-conquer strategy by partitioning extensive scenes into smaller sub-scenes based on camera projections or clustering of camera poses obtained from Structure-from-Motion (SfM), subsequently distributing these partitions across multiple GPUs [29, 20, 9, 33, 34]. While effective for controlled, uniform camera distributions such as five-directional aerial surveys [35, 54], these image-based methods struggle with close-range oblique photogrammetry scenarios, where irregular and complex camera trajectories result in uneven or inadequate training data distribution. Issue (2): Existing approaches necessitate merging sub-scenes before mesh extraction, a process requiring substantial computational resources. For example, methods employing TSDF fusion with a 0.2-meter voxel size consume over 300GB RAM for reconstructing merely 1,000 images, making scalability to city-scale scenes with tens of thousands of images impractical. Additionally, extracting meshes separately introduces further complications, such as determining optimal sub-scene boundaries, often leading to unnecessary computational overhead and fragmented meshes. Issue (3): Another significant challenge in GS-based large-scale reconstruction is the densification problem. GS explicitly represents scenes using discrete Gaussian primitives, meaning the richer and denser the primitives, the more accurately high-frequency geometric details can be captured [64, 10, 27, 13]. However, GS often struggles with accurately reconstructing low-frequency or textureless regions, resulting in incomplete or overly smooth surfaces [16, 13, 5]. Issue (4): Existing GS-based methods predominantly utilize single-view geometric consistency constraints during training [64, 10, 27, 13], which fail to leverage comprehensive multi-view information. This single-view approach can cause inconsistencies and inaccuracies, such as floaters or misaligned primitives, particularly in complex urban scenarios.
To address these critical issues, we propose ULSR-GS. Our key insight is combining detail-preserving optimization within each divided sub-region and enforcing multi-view consistency during training, which makes it possible for 2DGS to independently reconstruct each tile with high geometric fidelity. These sub-regions can then be seamlessly integrated into a globally consistent urban-scale surface without additional fusion or post-processing steps. Unlike other methods that partition scenes solely by camera positions [33, 67, 29, 34], we propose a point-to-photo partitioning strategy driven by the spatial distribution of sparse SfM points and their associated views. For each candidate sub-region, we meticulously select an optimal subset of images for each SfM point, ensuring comprehensive and relevant coverage. We then introduce a multi-view constrained densification approach within the optimization process for each partitioned sub-region. This strategy explicitly enforces geometric consistency across multiple views by penalizing discrepancies in depth and normal orientations of corresponding points across different images. Consequently, our multi-view densification preserves essential fine details, which are preserved when sub-regions are optimized independently. Collectively, the innovative integration of our point-to-photo partitioning and multi-view densification strategies positions ULSR-GS as a highly effective and scalable solution for accurate and detailed large-scale urban surface reconstructions.
@@
  • We present ULSR-GS, a novel method specifically designed to overcome the limitations of existing GS-based works in large-scale surface reconstruction tasks.
  • A partitioning strategy that supports extracting the geometry of each trained GS sub-region separately, without the need for merging and then extracting.
  • An densification method based on multi-view depth aggregation, which can improve rendering and geometry reconstruction quality.
  • Two efficient multi-view consistency constraints for generating detailed and accurate reconstructions in large-scale urban environments.
@@ 2.1 Large-Scale Novel View Synthesis
3D reconstruction has evolved from traditional Multi-View Stereo (MVS) [44, 24, 56, 57, 3, 59, 71, 41, 32, 60, 14, 31, 70, 28] to advanced neural methods [49, 40, 63], and mesh extraction methods transform implicit representations into explicit 3D models. Poisson Surface-based methods extract surfaces from point clouds but often produce overly smooth results, lacking fine details [15]. Marching Cubes converts volumetric data into polygonal meshes, but its grid-based nature can lead to artifacts and increased memory usage [36, 27]. Neural implicit representations encode 3D geometry continuously, enabling smooth modeling and handling complex topologies [49, 63, 62]. However, these methods require significant computational resources and often struggle with fine geometric details. Mesh extraction from these representations also suffers from discretization artifacts. Neural Radiance Fields (NeRF) use volumetric rendering to achieve high-quality view synthesis [39]. Despite their impressive results, NeRF-based methods are computationally intensive, making them less practical for real-time applications [1, 47]. Recent works extract explicit geometry from NeRF, but converting density fields to surfaces in large-scale scenes remains challenging and may produce noisy results [26].
@@
Figure 1: ULSR-GS training pipeline. Stage 1: Scene partitioning based on point cloud. Stage 2: Training strategy based on multi-view geometric consistency. Stage 3: Extract the geometry of sub-regions separately and seamlessly merge.
@@ 3.1 Multi-View Pair Selection
The function $G$ is a piecewise Gaussian function that favors a certain baseline angle $\theta_{0}$:
@@ 3.2 Depth Calculation in 2D Gaussian Splatting
$D_{mean}=\sum_{i}{\omega_{i}D_{i}}/(\sum_{i}{\omega_{i}}+\epsilon)$ (3)
@@
Maximum visibility. This method focuses on selecting the depth of the most "visible" Gaussian along the viewing ray. The depth $D_{median}$ is determined by the first Gaussian that contributes significantly to the cumulative alpha:
@@
$D_{mean}$ facilitates smoother depth calculations, whereas $D_{median}$ more effectively captures rich geometric detail. However, it is highly susceptible to floating-point errors that can impact depth computation. In this paper, we adopt $D_{mean}$ as the depth representation for subsequent analyses.
@@ 4 Method
Figure 1 illustrates the overall pipeline of our proposed ULSR-GS, which consists of three stages: Stage 1: Multi-View Optimized Point Partitioning (Sec. 4.1). Starting from a full-scene COLMAP SfM result, we apply density-aware boundary refinement (Sec. 4.1.1) followed by spatial subdivision of the sparse point cloud into multiple rectangular sub-regions (Sec. 4.1.2). Each point is then assigned to its most informative image group via our proposed per-point optimal view selection, resulting in a well-structured sub-region configuration suitable for parallel processing (Sec. 4.1.4). Stage 2: Sub-region Training. Each sub-region is independently trained using 2DGS rasterization. Our training module includes an adaptive multi-view guided densification mechanism (Sec. 4.2), where source view depths are aggregated (Sec. 4.2.1) and reprojected to the reference view through an adapted window mask (Sec. 4.2.2). To enforce geometric consistency across depth and normal maps, we further introduce two multi-view regulations (Sec. 4.3). This enables accurate local optimization with fine-grained geometry reconstruction. Stage 3: Sub-region Geometry Extraction. After training, surface meshes for each sub-region are extracted using TSDF fusion followed by a cropping step. These independently reconstructed regions are then seamlessly assembled into a complete large-scale urban mesh without requiring additional global fusion.
Figure 2: Point-based partition strategy of ULSR-GS. (I) We partition the training scene based on the positions of the point cloud projected onto the xz-plane. (II) For each sub-scene: (a) we first select all matching images within the region; (b) based on the camera relationships of the sub-scene, we select the corresponding images for each point; (c) following the principle of multi-view optimal matching [61], we filter out images with poor view angles; and finally, (d) considering the distance relationship, we further remove images that are too distant. (III) Examples of a sub-region significantly reduce the number of training images.
@@ 4.1 Multi-View Optimized Point Partitioning
In this section, we will detail the point-based scene partitioning method in Figure 1 Stage 1 and the method to select views for the partitioned sub-regions. Different from previous studies that use drone photo locations for region partitioning [29, 67, 33], our scene partitioning method is based on the initial point cloud of the scene and selects the best training images for the point cloud in each sub-region. The advantage of this method is that we can determine the boundaries of the mesh to be extracted for each sub-area at the early stage, without the need to merge the full scene. As illustrated in Figure 2, we optimize the partition process by focusing on the most relevant parts of the scene.
4.1.1 Density-controlled boundary refinement.
To effectively partition the initial points into $n\times n$ sub-regions, it is crucial to isolate the primary structural components of the scene and eliminate sparse and noisy SfM points that could distort the boundary definition.
@@
4.1.2 Initial view selection.
As shown in Figure 2 (I), the input COLMAP SfM [44] point cloud is divided into $n\times n$ grids after density filtering. Each point in the sub-region is the feature of the images in which it was detected and matched. We first select all matched images as a coarse view selection (see Figure 2 (e)).
4.1.3 Source view selection.
To further refine the view selection in Eq. (1), a region constraint is applied based on the camera pair distance $d_{ij}$. Only image pairs with a distance below a specified maximum threshold $D_{\max}$ are considered for matching. The final matching score $S_{ij}$:
@@
4.1.4 Per-point optimal view selection.
Our goal is to ensure that each SfM point within a sub-region is associated with its most informative and geometrically robust image pairs. For each point $P_{k}\in\mathcal{P}$, we consider all possible groups of images that observe $P_{k}$. Specifically, each group $\mathcal{G}_{i}={I_{\text{ref}_{i}},I_{\text{src}_{i}^{1}},I_{\text{src}_{i}^{2}},I_{\text{src}_{i}^{3}}}$ consists of one reference image and three corresponding source images. Our method involves the basic W2C steps.
We project $P_{k}$ onto the 2D image planes of each image $I_{j}\in\mathcal{G}_{i}$:
$d_{k_{j}}=\sqrt{(u_{k_{j}}-c_{x_{j}})^{2}+(v_{k_{j}}-c_{y_{j}})^{2}}.$ (10)
Among all possible four-image groups that observe $P_{k}$, we select the group with the smallest $D_{avg}(P_{k})_{min}$. This ensures that $P_{k}$ is primarily reconstructed using images where it is closest to the image centers in their respective 2D planes, thereby enhancing the reliability of its triangulation. Since we want to select the optimal image group for every initial SfM point in the sub-regions. We simply identify the group for $P_{k}$ $\mathcal{G}$ by minimizing $D_{avg}$:
$I_{used}=\bigcup_{P_{k}\in\mathcal{P}}\mathcal{G}_{opt}(P_{k})$ (13)
@@
4.1.5 Scene Partitioning Strategy
To address the issue where some sub-regions may contain too few points (especially in boundary areas) after dividing the scene into smaller parts, we evaluate each sub-region as follows: We retain a sub-region only if both the number of points and the number of matched images within it reach at least 10 % of the average per-subregion counts (i.e., the total number of points or images divided by $n^{2}$).
Table 1: Example Per Scene Partitioning Strategy.
Scene Images Grid Size Used
Campus [35] 5871 $8\times 8$ 41
Residence [35] 2582 $6\times 6$ 20
SZTU [54] 1500 $6\times 6$ 17
LOWER_CAMP [54] 670 $4\times 4$ 12
We partition each scene based on the number of images to ensure efficient processing and optimal reconstruction quality. Specifically, we divide the regions according to the following criteria:
  1. 1.
    Scenes with fewer than 1,000 images are divided into a $4\times 4$ grid.
  2. 2.
    Scenes with fewer than 3,000 images are divided into a $6\times 6$ grid.
  3. 3.
    Scenes with more than 3,000 images are divided into an $8\times 8$ grid.
@@ 4.2 Adaptive Multi-View Densification
In this section, we detail how multi-view geometric consistency is exploited to introduce densification based on rendered depth maps for 2DGS. As demonstrated in Figure 1 (Stage 2), each subregion can be assigned to an independent GPU for training. However, vanilla 2DGS often struggles to capture fine geometry details due to densification based on single-view image gradients, resulting in low-fidelity rendering and overly smooth geometry extraction. Therefore, it is necessary to add more fine Gaussian primitives for capturing high-frequency geometric details [13, 27, 10, 66]. In ULSR-GS, we perform additional densification via an MVS-like approach to address the limitations of overly smooth meshes resulting from $D_{mean}$ computations in Eq. 3 from TSDF fusion. This method projects $D_{mean}$ into 3D space [61, 11] and combines it with the RGB information of the GT image to enrich the Gaussian primitives.
4.2.1 Multi-view depth aggregation.
In our approach, we perform weighted averaging of depth from multiple source views. This weighting assigns a confidence score to each depth estimate based on its geometric consistency [61], ensuring that source views with higher consistency have a greater influence on the final depth estimation.
Since we have $3$ source views from Eq. (8), with each rendered depth $D_{mean}$ denoted as $\mathbf{D}_{\text{src}_{i}}$ for $i=1,2,3$, and the depth map of the reference view denoted as $\mathbf{D}_{\text{ref}}$. For each pixel $p_{\text{ref}}$ in the reference view, the final fused depth estimate $\mathbf{D}_{\text{final}}(p_{\text{ref}})$ is obtained through weighted fusion of the source views:
@@
Table 2: Quantitative results on GauU-Scene dataset [54]. We report precision, recall, and F-1 score at the threshold $\tau=0.025$ (relative). The best and second results are highlighted. "OOM" denotes no results due to GPU out of memory. " *" denotes no result due to no convergence.
Scene LOWER_CAMP UPPER_CAMP LFLS SMBU SZIIT SZTU
Neuralangelo * * * * * * * * * * * * * * * * * * $>$100 h
GOF-60K 0.667 0.652 0.659 0.693 0.672 0.682 0.607 0.659 0.631 0.679 0.784 0.727 0.649 0.695 0.671 OOM OOM OOM 5.3h
PGSR-60K 0.683 0.648 0.665 0.672 0.673 0.672 OOM OOM OOM 0.660 0.795 0.721 0.679 0.642 0.659 0.548 0.542 0.545 5.7h
Ours 0.681 0.689 0.685 0.690 0.705 0.697 0.684 0.703 0.693 0.737 0.789 0.762 0.705 0.713 0.709 0.713 0.723 0.718 3.7h
Table 3: Quantitative results on Matrix City dataset (Small City Scene) [25] and two custom-collected datasets (Scene1 and Scene2). We report precision, recall, and F-1 score at different distance thresholds $\tau=1.0m$ and $\tau=0.5m$ (metric). The best, second results are highlighted. "OOM" denotes no results due to GPU out of memory. Note that "City-GS V2+Ours(2$\times$2)" means we train CityGaussianV2 [34] on our proposed partition method.
Matrix City Scene1 Scene2
$\tau=1.0m$ $\tau=0.5m$ $\tau=1.0m$ $\tau=0.5m$ $\tau=1.0m$ $\tau=0.5m$
Metrics Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$ Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$ Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$ Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$ Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$ Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$
2DGS-60K (Baseline) 0.551 0.692 0.613 0.331 0.502 0.398 OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
City-GS V2 0.871 0.901 0.886 0.667 0.753 0.707 OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
City-GS V2+Ours(2$\times$2) - - - - - - 0.845 0.875 0.860 0.646 0.753 0.695 0.702 0.746 0.723 0.569 0.637 0.601
Ours 0.869 0.916 0.892 0.660 0.759 0.706 0.837 0.929 0.880 0.632 0.787 0.701 0.710 0.748 0.728 0.569 0.651 0.607
4.2.2 Adaptive depth densification.
Directly projecting all depth as new 2D Gaussian primitives after checking the geometric consistency introduces excessive redundant information into the training scene [27]. This redundancy negatively impacts both the training speed and the accuracy of the scene representation. To address this issue, we introduce an adaptive densification window mask that confines the densification area. This approach eliminates inaccurate values at the edges of the depth map and adaptively accounts for non-uniform depth changes caused by varying viewpoint poses.